Find all solutions of the given differential equations and then find the particular solutions for which a point (x,y) is given:


dy/dx = -32x + 8 ; (x,y) = (0,1)


1
Expert's answer
2021-09-03T02:40:59-0400
"\dfrac{dy}{dx}=-32x+8"

"dy=(-32x+8)dx"

Integrate


"\int dy=\int(-32x+8)dx"

"y=-16x^2+8x+C"

The general solution of the given differential equation is


"y(x)=-16x^2+8x+C"



Point "(x,y)=(0,1)"


"1=-16(0)^2+8(0)+C=>C=1"

The particular solution is


"y=-16x^2+8x+1"